OpenStudy (anonymous):

3x-y=10-x+y=-4 3,-7 -1,3 3,-1 3,1

OpenStudy (anonymous):

hey man, i wrote this wrong, i put this another time :) check onthe other side

OpenStudy (anonymous):

3x-x=10-4 2x=6 x=3, y=0